# DiBS-MTL: Transformation-Invariant Multitask Learning with Direction Oracles

### v1

@@ 6 Conclusion
We consider the problem of constructing multitask learning (MTL) methods that are robust to monotonic, nonaffine task loss transformations. In practice, different task losses can be arbitrarily scaled with respect to one another, and those scalings can be understood as monotonic, possibly nonaffine transformations of some ideal, unknown losses which are meaningfully comparable. To address this problem, we present DiBS-MTL, an MTL method which is invariant to such transformations because it uses only normalized task loss gradients. Building upon recent work which formalizes the MTL training update step as a bargaining game played between tasks (Navon et al., 2022), we adapt the recently introduced DiBS bargaining approach (Gupta et al., 2025) in order to find a Pareto stationary (and monotonic transformation-invariant) update direction. While DiBS enjoys convergence guarantees when losses are strongly convex, practical MTL problems almost always have nonconvex task losses. To bridge this gap, we prove that a subsequence of the DiBS iterates asymptotically converges to a Pareto stationary point even in the nonconvex regime. Empirically, we demonstrate that DiBS-MTL achieves performance competitive with state-of-the-art methods on standard large-scale MTL benchmarks. Moreover, when task losses are nonaffinely transformed other leading MTL methods suffer degraded performance, with some tasks dominating others, whereas DiBS-MTL still maintains high performance, demonstrating robustness to such transformations.
@@ Appendix C Discussion on IMTL-G
Appendix D Experimental Details and Code base
D.1 demonstrative Example
@@ D.2 NYuV2
The NYU-v2 indoor scene dataset Silberman et al. (2012) provides 1,449 RGB-D images with dense per-pixel annotations for three tasks: semantic segmentation (13 classes), depth estimation, and surface normal prediction. Training uses standard task losses: per-pixel cross-entropy for segmentation, masked L1 for depth estimation over valid pixels, and a masked cosine loss over valid pixels for surface normal prediction. These choices match the NYU-v2 multitask protocol followed by prior multitask learning work (Liu et al., 2023; Navon et al., 2022). Our implementation builds on the official Nash-MTL code base [^4] https://github.com/AvivNavon/nash-mtl and incorporates the FAMO implementation [^5] https://github.com/Cranial-XIX/FAMO.
@@
The model architecture used in our experiments is a Multitask Attention Network (MTAN) (Liu et al., 2019) built on top of SegNet (Badrinarayanan et al., 2017). The network takes an RGB image as input and produces three task-specific outputs via separate heads: (i) a per-pixel class-score map for semantic segmentation, (ii) a scalar depth map for depth estimation, and (iii) a 3-channel surface-normal map for surface-normal prediction. Following the experimental setup of Nash-MTL (Navon et al., 2022), all multitask learning methods train this shared architecture using their respective update rules.
@@
For training, we follow the procedure outlined in (Navon et al., 2022). We train for 200 epochs with Adam (Kingma & Ba, 2014), using an initial learning rate of $1\times 10^{-4}$ reduced to $5\times 10^{-5}$ after 100 epochs. All reported results are averaged over three random seeds, namely $1,7$, and $42$.
D.3 Meta-World MT10
Benchmark and Setup.
The MetaWorld MT10 (v1) benchmark comprises ten robotic manipulation tasks: Reach, Push, Pick-and-Place, Door Open, Drawer Open, Drawer Close, Button Press (Top-Down), Peg Insert (Side), Window Open, and Window Close (Yu et al., 2020b). Each task has distinct reward functions and success criteria. We build on the MTRL code base (Sodhani & Zhang, 2021) [^6] https://github.com/facebookresearch/mtrl with MetaWorld [^7] https://github.com/Farama-Foundation/Metaworld.
Policy and Training Protocol.
We train a single Soft Actor–Critic (SAC) policy shared across all ten tasks. Multitask learning (MTL) methods are applied to balance the actor and critic updates within the shared SAC. For each MTL method, we train for 2,000,0002{,}000{,}000 environment steps in total. Episodes have length $150$; this corresponds to $\approx 13{,}333$ episodes overall. Evaluation is performed every $200$ episodes (30,00030{,}000 steps) on all tasks, and reported metrics are averaged over 10 random seeds: $1,2,3,4,5,6,7,8,9,10$.
@@ D.3 Meta-World MT10
Key hyperparameter values are provided in Table 3. We follow the same hyperparemeters defined in prior works  (Navon et al., 2022; Liu et al., 2023).
@@
During preliminary experiments, we observed that using the same random seed did not always produce identical training curves. This discrepancy arose because the original MTRL code (Sodhani & Zhang, 2021) did not seed the initial action sampling during the exploration phase. As a result, small differences in early exploratory actions (first 10 steps) propagated through training and produced slight variation in learning behavior even with fixed seeds.
Appendix E Additional Results
E.1 Analysis of DiBS-MTL  Runtime
Nash-MTLDiBS-MTLFAMO01,0001{,}0002,0002{,}000108.32108.321,092.871{,}092.872,125.962{,}125.96Iterations per second
Figure 4: Average iterations per second in the non-transformed setting
In addition to examining the solutions obtained in the demonstrative example, we also evaluated the computational speed of the different MTL methods. Specifically, we compared DiBS-MTL  with Nash-MTL and FAMO. We measured the number of iterations processed per second, with higher values indicating faster performance.
In Figure 4, we compare the runtimes of DiBS-MTL , FAMO, and Nash-MTL. DiBS-MTL  runs substantially faster than Nash-MTL, achieving roughly half the speed of FAMO. This behavior is expected: unlike Nash-MTL, DiBS-MTL  does not solve a separate optimization problem at each iteration, but it does recompute gradients and preferred states, which increases runtime relative to FAMO, which is a loss-based method and is expected to be faster.
(a) Nominal
@@ E.1 Analysis of DiBS-MTL  Runtime
Figure 6: Additional experiments performed in the demonstrative two-objective example. We observe that, Multi-step DiBS-MTL  exhibits the same invariance to monotone non-affine transforms as single-step DiBS-MTL, and also demonstrates similar behaviour.
E.2 Additional Results in the demonstrative Example
In addition to the results reported in section Section 5.1, we also ran additional methods in the demonstrative two-objective example. Specifically, we ran the Multiple Gradient Descent Algorithm (MGDA)  (Désidéri, 2012), Uncertainty Weighting (UW)  (Kendall et al., 2018), Impartial Multitask Learning (IMTL-G)  (Liu et al., 2021b), and Multi-step DiBS-MTL. We run Multi-step DiBS-MTL for 10 steps per update. The loss functions and number of steps are identical to the experiment described in Section 5.1.
In Figure 6, we observe that the additional baseline methods of MGDA, UW, IMTL-G are not invariant to the transform, with UW failing to reach the Pareto-front in both cases. We note that Multi-step DiBS-MTL performs similarly to single step DiBS-MTL . Multi-step DiBS-MTL reaches the same points on the Pareto-front in both cases, showing it is also invariant to monotone non-affine transforms.

### v2

@@ 6 Conclusion
We consider the problem of constructing multi-task learning (MTL) methods that are robust to monotonic, nonaffine task loss transformations. In practice, different task losses can be arbitrarily scaled with respect to one another, and those scalings can be understood as monotonic, possibly nonaffine transformations of some ideal, unknown losses which are meaningfully comparable. To address this problem, we present DiBS-MTL, an MTL method which is invariant to such transformations because it uses only normalized task loss gradients. Building upon recent work which formalizes the MTL training update step as a bargaining game played between tasks (Navon et al., 2022), we adapt the recently introduced DiBS bargaining approach (Gupta et al., 2025) in order to find a Pareto stationary (and monotonic transformation-invariant) update direction. While DiBS enjoys convergence guarantees when losses are strongly convex, practical MTL problems almost always have nonconvex task losses. To bridge this gap, we prove that a subsequence of the DiBS iterates asymptotically converges to a Pareto stationary point even in the nonconvex regime. Empirically, we demonstrate that (i) due to its monotonic transformation invariance, DiBS-MTL significantly outperforms state-of-the-art MTL baselines in settings with task losses being poorly scaled relative to each other, highlighting the unique suitability of DiBS-MTL to MTL problems with heterogeneous, possibly unknown task loss scalings; and (ii) the transformation robustness property of DiBS-MTL does not sacrifice performance on real-world benchmarks, where DiBS-MTL is competitive against leading MTL baselines.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix C Discussion on IMTL-G
Appendix D Additional Supervised MTL Experiments – Cityscapes benchmark
Table 4: Results for supervised learning Cityscapes benchmark. Segmentation metrics (mIoU, Pix Acc) are shown in percentage form. Bold, underlined indicate best and second-best per column.
Segmentation (%) Depth Estimation
mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ $\Delta m\%$ $\downarrow$ MR $\downarrow$
STL 74.01 93.16 0.0125 27.77
MGDA 69.44 91.49 0.0137 45.26 20.03 6.3
UW 66.18 90.28 0.0158 54.92 34.41 10.2
Nash-MTL 73.57 92.92 0.0140 52.34 25.30 6.0
FAMO 69.80 91.66 0.0137 43.66 18.51 5.5
LS 71.89 92.44 0.0180 111.89 87.64 10.0
GO4ALIGN 72.38 92.60 0.0312 93.50 97.26 9.8
IGBV2 66.44 90.48 0.0153 56.56 34.85 10.0
IGBV2+Nash 73.53 92.95 0.0135 50.42 22.58 4.5
LDC 75.04 93.44 0.0144 38.44 12.93 3.0
DiBS-MTL 74.64 93.31 0.0135 58.09 29.10 5.9
DiBS-MTL (5 Steps) 75.46 93.55 0.0132 54.88 25.27 3.2
DiBS-MTL (10 Steps) 75.05 93.41 0.0133 52.36 23.34 3.6
The Cityscapes dataset (Cordts et al., 2016) contains 5000 street-view $128\times 256$, RGB images annotated for two tasks: (1) semantic segmentation with 19 classes and (2) discretized depth estimation with 7 ordinal bins. Following prior MTL evaluations on Cityscapes (Navon et al., 2022; Liu et al., 2023), we report mIoU and pixel accuracy for segmentation, and absolute and relative depth error for the depth task. This dataset complements NYU-v2 by providing a larger-scale, outdoor benchmark with visually and semantically different task relationships.
Results on Cityscapes.
In the Cityscapes benchmark, DiBS-MTL variants similarly perform competitively against all baselines, achieving strong segmentation performance and favorable depth metrics across the two tasks (see Table 4), and the 5 step variant of DiBS-MTL achieves the second best MR.
Appendix E Experimental Details and Code base
E.1 Demonstrative Nonconvex Example
@@ E.2 NYuV2
The NYU-v2 indoor scene dataset (Silberman et al., 2012) provides 1,449 RGB-D images with dense per-pixel annotations for three tasks: semantic segmentation (13 classes), depth estimation, and surface normal prediction. Training uses standard task losses: per-pixel cross-entropy for segmentation, masked L1 for depth estimation over valid pixels, and a masked cosine loss over valid pixels for surface normal prediction. These choices match the NYU-v2 multi-task protocol followed by prior multi-task learning work (Liu et al., 2023; Navon et al., 2022). Our implementation builds on the official Nash-MTL code base [^2] https://github.com/AvivNavon/nash-mtl and incorporates the FAMO implementation [^3] https://github.com/Cranial-XIX/FAMO. Following standard practice (Liu et al., 2019; Navon et al., 2022; Liu et al., 2023; Xiao et al., 2025; Shen et al., 2024; Dai et al., 2023), we evaluate segmentation using mean Intersection-over-Union (mIoU) and pixel accuracy; depth estimation using mean absolute error and mean relative error; and surface normals using median angular error, mean angular error, and the percentage of pixels whose angular error is below $30^{\circ}$, $22.5^{\circ}$, and $11.25^{\circ}$ (Silberman et al., 2012). These metrics provide a comprehensive view of per-task performance across classification, regression, and geometric prediction settings.
@@
The model architecture used in our experiments is a Multi-Task Attention Network (MTAN) (Liu et al., 2019) built on top of SegNet (Badrinarayanan et al., 2017). The network takes an RGB image as input and produces three task-specific outputs via separate heads: (i) a per-pixel class-score map for semantic segmentation, (ii) a scalar depth map for depth estimation, and (iii) a 3-channel surface-normal map for surface-normal prediction. For LS, MGDA, UW, FAMO, NashMTL, IGBv2, and DiBS-MTL, we use the multi-task attention network (MTAN) (Liu et al., 2019) with a SegNet-style encoder–decoder architecture and apply the same training schedule, optimizer, and data-augmentation procedures as established in prior work for a fair comparison. For LDC (Xiao et al., 2025) and GO4Align (Shen et al., 2024), we use the task architectures provided in their original implementations.
@@
For training, we follow the procedure outlined in (Navon et al., 2022). We train for 200 epochs with Adam (Kingma and Ba, 2014), using an initial learning rate of $1\times 10^{-4}$ reduced to $5\times 10^{-5}$ after 100 epochs. All reported results are averaged over three random seeds, namely $1,7$, and $42$.
@@
E.3 Cityscapes
Benchmark and Setup.
The Cityscapes outdoor driving scenes (Cordts et al., 2016) provides 5000 RGB-D images with dense per-pixel annotations for 2 tasks: semantic segmentation (7 classes) and depth estimation. Training uses standard task losses: per-pixel cross-entropy for segmentation and masked L1 for depth estimation over valid pixels These choices match the Cityscapes multi-task protocol followed by prior multi-task learning work (Liu et al., 2023; Navon et al., 2022).
Model.
Just as with NYU-v2, the model architecture used in our experiments is a Multi-Task Attention Network (MTAN) (Liu et al., 2019) built on top of SegNet (Badrinarayanan et al., 2017). The network takes an RGB image as input and produces two task-specific outputs via separate heads: (i) a per-pixel class-score map for semantic segmentation, (ii) a scalar depth map for depth estimation. For LS, MGDA, UW, FAMO, NashMTL, IGBv2, and DiBS-MTL, we use the multi-task attention network (MTAN) (Liu et al., 2019) with a SegNet-style encoder–decoder architecture and apply the same training schedule, optimizer, and data-augmentation procedures as established in prior work for a fair comparison. For LDC (Xiao et al., 2025) and GO4Align (Shen et al., 2024), we use the task architectures provided in their original implementations.
Training Protocol.
We follow the procedure outlined in (Navon et al., 2022). We train for 300 epochs with Adam (Kingma and Ba, 2014), using an initial learning rate of $1\times 10^{-4}$ reduced to $5\times 10^{-5}$ after 100 epochs. All reported results are averaged over three random seeds, namely $1,7$, and $42$.
E.4 QM9
Benchmark and Setup.
The QM9 dataset (Blum and Reymond, 2009) is a graph neural network learning becnhmark consisting of 11 tasks, and is a widely used challenging benchmark in MTL literature (Navon et al., 2022; Liu et al., 2023; Xiao et al., 2025; Dai et al., 2023; Shen et al., 2024). It consists of molecules represented by annotated graphs. Our experimental setting is the same as that used in prior MTL works, and the tasks are to predict 11 different molecular properties.
Model.
Following Navon et al. (2022), we use the PyTorch Geometric implementation of Fey and Lenssen (2019), which incorporates the widely used GNN architecture of Gilmer et al. (2017) together with the pooling operator of Vinyals et al. (2015). As with the other benchmarks, LDC and GO4Align use their own architectures.
Training Protocol.
We train each method for 300 epochs and perform a learning-rate search based on the validation $\Delta m$ metric. A learning-rate scheduler reduces the learning rate once the validation $\Delta m$ stops improving, and the validation set is additionally used for early stopping. All reported results are averaged over three random seeds, namely 1, 7, and 42. This is in line with the approach taken by  (Navon et al., 2022).
E.5 Meta-World MT10
Benchmark and Setup.
The MetaWorld MT10 (v1) benchmark comprises ten robotic manipulation tasks: Reach, Push, Pick-and-Place, Door Open, Drawer Open, Drawer Close, Button Press (Top-Down), Peg Insert , Window Open, and Window Close (Yu et al., 2020b). Each task has distinct reward functions and success criteria. We build on the MTRL code base (Sodhani and Zhang, 2021) [^4] https://github.com/facebookresearch/mtrl with MetaWorld [^5] https://github.com/Farama-Foundation/Metaworld.
Policy and Training Protocol.
We train a single Soft Actor–Critic (SAC) policy shared across all ten tasks. multi-task learning (MTL) methods are applied to balance the actor and critic updates within the shared SAC. For each MTL method, we train for 2,000,0002{,}000{,}000 environment steps in total. Episodes have length $150$; this corresponds to $\approx 13{,}333$ episodes overall. Evaluation is performed every $200$ episodes (30,00030{,}000 steps) on all tasks, and reported metrics are averaged over 10 random seeds: $1,2,3,4,5,6,7,8,9,10$.
[…]
(a) Average seconds per iteration ($\downarrow$ is faster) in the nonconvex example from Section 5.1.
$Nash-MTL$ 1 in 100 steps1-Step DiBS-MTL020204040606049.9449.948.218.216.336.33Seconds per epoch
(b) Average seconds per epoch ($\downarrow$ is faster) for the multi-task RL example in Section 5.2.
Figure 4: Runtime comparisons across nonconvex demonstrative example and multi-task RL settings.
Table 6: Run times in minutes ($\downarrow$ is faster) for supervised MTL experiments in Section 5.3 averaged over 3 seeds.
Benchmarks
NYU-V2 Cityscapes QM9
1-step DiBS-MTL 819.33 481.67 1705
5-step DiBS-MTL 1713 612 1719
10-step DiBS-MTL 1799 752 1726
Nash-MTL 1031 598 1699
@@ E.5 Meta-World MT10
Key hyperparameter values are provided in Table 5. We follow the same hyperparemeters defined in prior works  (Navon et al., 2022; Liu et al., 2023, 2021a; Ban and Ji, 2024).
@@
During preliminary experiments, we observed that using the same random seed did not always produce identical training curves. This discrepancy arose because the original MTRL code (Sodhani and Zhang, 2021) did not seed the initial action sampling during the exploration phase. As a result, small differences in early exploratory actions (first 10 steps) propagated through training and produced slight variation in learning behavior even with fixed seeds.
Appendix F Additional Results
F.1 Comparison of DiBS-MTL  runtimes with previous bargaining MTL method Nash-MTL.
We also evaluated the computational speed of the DiBS-MTL with the previous bargaining-inspired MTL method, Nash-MTL. Specifically, we compare runtimes of DiBS-MTL  with Nash-MTL for the demonstrative nonconvex examples (Figure 4), the multi-task reinforcement learning example (Figure 4), and the supervised learning NYU-v2, QM9 and Cityscapes examples (Table 6).
From Figure 4, it is clear that DiBS-MTL consistently is faster than Nash-MTL for the demonstrative nonconvex and the multi-task reinforcement learning examples. In the multi-task reinforcement example particularly, Nash-MTl is particularly slow if applied for every training iteration, and even when applied only once per 100 training loops (which is the speedup suggested by the original authors (Navon et al., 2022)), it is still slower than 1-step DiBS-MTL. We conjecture that the slow performance of Nash-MTL is because it solves a separate convex optimization problem at each iteration which is time consuming.
In Table 6, we compare the runtime of different DiBS-MTL  variants and Nash-MTL across the MTL benchmarks to assess how the number of bargaining steps affects computational cost. The single-step version of DiBS-MTL  achieves the fastest runtimes on NYU-V2 and Cityscapes, while falling slightly behind Nash-MTL on QM9. As expected, runtime for DiBS-MTL variants increases as the number of steps increase, because the per epoch complexity increases, reflecting the added cost of performing multiple bargaining updates. At the same time, as indicated by Tables 2, 3 and 4, 1-step DiBS-MTL achieves comparable performance to the multi-step versions, making 1-step DiBS-MTL an appropriate approximation of multi-step DiBS-MTL variants.
(a) Untransformed
@@ Appendix F Additional Results
Figure 6: Additional experiments performed in the demonstrative two-objective example. We observe that, 1-step DiBS-MTL exhibits the same invariance to monotone non-affine transforms as T-step DiBS-MTL from Figure 2, and also demonstrates similar behaviour, tracing very similar trajectories as T-step DiBS-MTL ($T=5$) from the same initializations.
Algorithm 2 DiBS-MTL Single Step
0: Initial parameters $\theta_{(0)}$, losses $\{\ell^{i}\}_{i=1}^{N}$, learning rate $\eta$, number of epochs $J$
for $j=1$ to $J$ do
  Compute task gradients $\nabla_{\theta_{(j-1)}}\ell^{i}$
  Compute DiBS update direction: $d_{(j)}=\sum_{i\in[N]}\frac{\nabla_{\theta}\ell^{i}(\theta_{(j-1)})}{\|\nabla_{\theta}\ell^{i}(\theta_{(j-1)})\|_{2}}$
  Update: $\theta_{(j)}\leftarrow\theta_{(j-1)}-\eta\,d_{(j)}$
end for
return $\theta_{(J)}$
F.2 Additional Results in the Demonstrative Nonconvex Example
In addition to the results reported in section Section 5.1, we also ran additional methods in the demonstrative two-objective example. Specifically, we ran the Multiple Gradient Descent Algorithm (MGDA)  (Désidéri, 2012), Uncertainty Weighting (UW)  (Kendall et al., 2018), Impartial Multi-task Learning (IMTL-G)  (Liu et al., 2021b), and T-step DiBS-MTL. We run T-step DiBS-MTL for 10 steps per update. The loss functions and number of steps are identical to the experiment described in Section 5.1.
In Figure 6, we observe that the additional baseline methods of MGDA, UW, IMTL-G are not invariant to the transform, with UW failing to reach the Pareto-front in both cases. We note that T-step DiBS-MTL performs similarly to single step DiBS-MTL . T-step DiBS-MTL reaches the same points on the Pareto-front in both cases, showing it is also invariant to monotone non-affine transforms.
Appendix G 1-step DiBS-MTL Algorithm Pseudocode
Algorithm 2 provides pseudocode for the 1-step DiBS-MTL algorithm. When adapting DiBS-MTL  to the single-step variant, the distances to the preferred states for each agent are equal. For 1-step DiBS-MTL, for all $j$ in Algorithm 1, $\|\theta_{(j-1)}-\theta^{*,i}\|=\epsilon$ for all $i$. The approximation radius $\epsilon$ and the step size $\eta_{s}$ are absorbed into the learning rate $\eta$.